### v1

# On Sparsity of integral points in orbits with big pullbacks under iterates

Abstract.
We prove new unconditional results of sparsity of integral points on orbits under many maps in arbitrary dimensions, generalizing theorems of Yasufuku(2015) and others. The main ingredients are new diophantine approximation tools.
1. Introduction
Let $\phi:X\rightarrow X$ be a self rational map on a projective variety defined over a number field $k$ and denote by $\phi^{(n)}$ the $n$-iterate of $\phi$. We investigate sparsity in the currently very active field of study about the frequency of integral points lying in the forward orbits
$\mathcal{O}_{\phi}(P):=\{P,\phi(P),\phi^{(2)}(P),...\}$
with $P\in X(k)$ in arithmetic dynamics, particularly in higher dimensions. In dimension $1$, finiteness of algebraic integers in orbits of rational functions was proven in a key breakthrough result of Silverman [16], who showed the only exceptions are maps whose second iterate is a polynomial. This was made quantitative and extended to "quasiintegral points" [8], to semigroup dynamics [15, 16], and also consistently described over function fields in [2, 9, 14] recently, just to cite a few. As in the classical cases of curves and abelian varieties, the main tools to show sparsity of integral points have been mostly powerful diophantine approximation results like Roth’s theorem or Schmidt’s subspace theorem, although in higher dimensions the problem is much more complicated. Zariski non-density of integral points in varieties of general type should be true according the Bombieri-Lang conjecture. This is also a consequence of a deep diophantine approximation conjecture due to Vojta that generalizes Roth and Schmidt, and also implies the ABC conjecture. Some of the first generalizations of Silverman’s theorem to higher dimensions are conditional to Vojta’s conjecture, as in the following nice sparsity theorem of Yasufuku on projective spaces
Theorem 1.1.
(Yasufuku)[20] Let $\phi:\mathbb{P}^{N}\rightarrow\mathbb{P}^{N}$ be a morphism defined over $\mathbb{Q}$ of degree $d\geq 2$ and $D$ be a divisor of $\mathbb{P}^{N}$. Suppose there exists $m>0$ such that the normal-crossings subdivisor $D_{nc}^{(m)}$ of the pullback $(\phi^{(m)})^{*}(D)$ satisfies Vojta’s conjecture and $\deg D_{nc}^{(m)}>N+1$. Then, for any $P\in\mathbb{P}^{N}(\mathbb{Q})$, the set $\mathcal{O}_{\phi}(P)\cap\mathbb{P}^{N}(\mathbb{Z}\setminus|D|)$ is not Zariski dense.
The theorem holds more generally over a number field $k$ and replacing integral points with the set of $(D,S)$-integral points
$\left\{P\in\mathbb{P}^{N}(k)|\sum_{v\notin S}\lambda_{v}(D,P)\leq ch(D,P)\right\}$
where $S$ is a finite set of places of $k$, $c$ is a positive constant and $\lambda_{v}(D,.)$ is a Weil local height function and $h(D,.)$ is a global Weil height that we recall further below. It also agrees with Silverman’s result in one variable. Other recent higher dimensional dynamical analogues of Siegel’s theorem for curves like the previous one can be found in recent works [3, 4, 13, 20, 21], either relying on the conjectures of Vojta, or on new generalizations of the Schmidt’s subspace theorem or an effective Runge-type result. Here, among other things we start using a new generalization of Schmidt’s subspace theorem due to He and Ru [7] to obtain an unconditional version of Yasufuku’s Theorem for projective varieties as follows
Theorem 1.2.
(See Theorem 4.1) Let $X$ be a projective variety of dimension $n$, $S$ be a finite set of places and $\phi$ be an endomorphism of $X$. Suppose there exist $m\geq 0$, a nontrivial effective divisor $D$ and $A$ an ample divisor of $X$ such that
  • (i)
    $(\phi^{(m)})^{*}(D)=D_{1}+D_{2}+\cdots+D_{q}+D^{\prime}$ is a sum of effective divisors where $D_{1},\ldots,D_{q}$ are in $l$-subgeneral position for some $l\geq n$
  • (ii)
    There exists $c_{j}\in\mathbb{Q}$ such that $\sum_{j=1}^{q}c_{j}{D_{j}}-(l-n+1)(n+1)A$ is big and $A-c_{j}D_{j}$ is $\mathbb{Q}$-nef for all $j\leq m$
Then there exists $\epsilon>0$ such that for all $P\in X(k)$, the set
$\big{\{}\phi^{(n)}(P):n\in\mathbb{N},\sum_{v\notin S}\lambda_{v}(D,\phi^{(n)}(P)\leq\epsilon h(A,\phi^{(n)}(P))\big{\}}$
is not Zariski dense.
This theorem recovers and generalizes Yasufuku’s results when his $D_{(nc)}^{m}$ is a union of lines and his theorem becomes unconditional due to Schmidt’s subspace theorem. Yasufuku’s theorem can be made quantitative with respect to bounding the largest number of iterations occurring in an integral iterate (even if we allow the base field to vary with bounded degree) and also uniform not depending on the initial point of the orbit. This was recently achieved by Yasufuku in [21] using new diophantine approximation ingredients. His uniformity result replacing points in a given number field with algebraic points of bounded degree provides a version of the univariate results of Gunther and Hindes [5] for $\mathbb{P}^{N}.$ Similarly, we prove a non-quantitative algebraic points analogue result for projective varieties assuming bigness of some iterates’ pullbacks of a base divisor as suggested in [21, Remark 8]. For dynamics with one map, we state that below
Theorem 1.3.
(See Theorem 4.4) Let $X$ be a projective variety of dimension $n$, $S$ be a finite set of places, and $\phi$ be an endomorphism of $X$. Suppose there exist $m$, a nontrivial effective divisor $D$ and $A$ an ample divisor of $X$ such that
  • (i)
    $(\phi^{(m)})^{*}(D)=D_{1}+D_{2}+\cdots+D_{q}+D^{\prime}$ a sum of effective ample divisors, $D_{1},\ldots,D_{q}$ are in $l$-subgeneral position with $l\geq n$,
  • (ii)
    There exist integers $d_{i}>0$ such that $\sum_{j=1}^{q}\frac{D_{j}}{d_{j}}-C(l,n,\delta)A$ is big for all $i$ and $D_{j}\equiv d_{j}A$ for all $j$ where $C(l,n,\delta)$ is a positive constant given in [11].
Then there exists $\epsilon>0$ such that the set
$\big{\{}\phi^{(n)}(P):n\in\mathbb{N}^{\geq m},[k(P):k]\leq\delta,\sum_{v\notin S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}\lambda_{w}(D,\phi^{(n)}(P)\leq\epsilon h(A,\phi^{(n)}(P)\big{\}}$
is not Zariski dense.
If $X=\mathbb{P}^{2}$ and the $D_{j}$’s are lines in general position and $A\sim O_{\mathbb{P}^{2}}(1)$, then condition (ii) can be reduced to $q>15/2$ thanks to a sharp diophantine approximation result by Levin [12, Theorem 1.5] (See Corollary 4.6).
Both theorems 1.2 and 1.3 are first proven and stated in the more general context of semigroup dynamics generated by finitely many maps(non-commuting in general).
In Section 2 we discuss some preliminaries on heights. The diophantine approximation tools are stated in Section 3, and the main results are proven in Sections 4 for nef and big pullbacks using He and Ru’s theorem and Le’s and Levin’s theorem on algebraic points respectively.
2. Preliminaries on heights
Let $k$ be a number field, and let $M_{k}$ be the set of places. For the unique archimedean place of $M_{\mathbb{Q}}$, we use the usual absolute value on $\mathbb{R}$ as the normalized one. For the non-archimedean place of $M_{\mathbb{Q}}$ corresponding to the prime $p$, we normalize the absolute value by defining the absolute value of $p$ to be $\dfrac{1}{p}$. We then normalize absolute value corresponding to each place $v\in M_{k}$ by defining $|x|_{v}$ to be the $\frac{[k_{v}:\mathbb{Q}_{v}]}{[k:\mathbb{Q}]}$-th power of the absolute value in $v$ whose restriction to $\mathbb{Q}$ is a normalized absolute value on $\mathbb{Q}$. We then define a Weil (global) height on the projective space $\mathbb{P}^{N}$ by
$h([a_{0}:\cdots:a_{N}])=\sum_{v\in M_{k}}\log\max_{i}|a_{i}|_{v}$
for $[a_{0}:\cdots:a_{N}]\in\mathbb{P}^{N}(k)$. This becomes a well-defined function on $\mathbb{P}^{N}(\overline{\mathbb{Q}})$. For a nontrivial effective divisor $D$ on $\mathbb{P}^{N}$ defined over $k$, we choose a homogeneous polynomial $F$ of degree $d$ with coefficients in the ring $\mathcal{O}_{k}$ of integers of $k$ so that $D$ is defined by $F$, and we define a local height function for each $v\in M_{k}$ by
$\lambda_{v}(D,[a_{0}:\cdots:a_{N}])=\log\frac{(\max|a_{i}|_{v})^{d}}{|F(a_{0},\ldots,a_{N})|_{v}}.$
This is a well-defined function on $(X\setminus|D|)(k)$. For a general divisor $D$ on a projective variety $X$, we first write $D$ as a difference of two very ample Cartier divisors $D_{1}$ and $D_{2}$, where $D_{i}=\varphi^{*}(H)$ for some closed immersion $\varphi_{i}:X\longrightarrow\mathbb{P}^{N_{i}}$ and $H_{i}$ is a hyperplane in $\mathbb{P}^{N_{i}}$. We then define Weil height to be
$h(D,P)=h(\varphi_{1}(P))-h(\varphi_{2}(P)).$
Moreover, letting $s_{i,1},\ldots,s_{i,\ell_{i}}$ be a basis of global sections of the line bundle $\mathscr{L}(D_{i})$ and choosing one rational section $s$ of $\mathscr{L}(D)$, we can define local height to be
$\lambda_{v}(D,P)=\max_{m}\min_{\ell}\log\left|\big{(}s_{1,m}\otimes(s_{2,\ell}\otimes s)^{-1}\big{)}(P)\right|_{v},$
where we evaluate at $P$ via the isomorphism $\mathscr{L}(D_{1})\otimes(\mathscr{L}(D_{2})\otimes\mathscr{L}(D))^{-1}\cong\mathscr{O}_{X}$. This notion is well-defined in the sense that if we choose different parameters for the ample divisors or the global/rational sections, the two local height functions agree on all but finitely many places and even at those finitely many places, their difference is a bounded function. The local and Weil height are functorial with respect to pullbacks, namely,
$h(\varphi^{*}D,P)=h(D,\varphi(P))+O(1)\qquad\text{ and }\qquad\lambda_{v}(\varphi^{*}D,P)=\lambda_{v}(D,\varphi(P))+O(1)$
whenever $\varphi$ is a morphism of algebraic varieties, where equality is interpreted to mean agreement up to a bounded function (or up to bounded functions in just finitely many places).
When $D$ and $E$ are divisors of $X$, the following additive properties also hold for every $P\in X(k)$ where the functions are defined
$h(D+E,P)=h(D,P)+h(E,P)+O(1)\quad\text{ and }\quad\lambda_{v}(D+E,P)=\lambda_{v}(D,P)+\lambda_{v}(E,P)+O(1)$
Also, if $D$ is effective, then
$h(D,P)\geq O(1)\quad\text{ and }\quad\lambda_{v}(D,P)\geq O(1)$
for all $P\in(X\setminus|D|)(k)$.
Moreover, it follows from the definitions that
(2.1) $h(D,P)=\sum_{v\in M_{k}}\lambda_{v}(D,P)+O(1)$
for all $P\in(X\setminus|D|)(k)$
3. Tools from diophantine approximation
Here we state the diophantine approximation results we will use. First, the following recent by He and Ru. For the definitions of general position and subgeneral position, see [7, 12]
Lemma 3.1.
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$ Let $S$ be a finite set of places of $k$. Let $D_{1},...,D_{q}$ be effective Cartier divisors on $X$, defined over $k$, in $l$-subgeneral position for $l\geq n$. Let $A$ be an ample Cartier divisor on $X$ and let $c_{i}$ be rational numbers such that $A-c_{i}D_{i}$ is a nef $\mathbb{Q}$-divisor for all $i$. Let $\epsilon>0$. Then there exists a proper Zariski-closed subset $Z\subset X$such that for all points $P\in X(k)\setminus Z$,
$\sum_{v\in S}\sum_{i=1}^{q}c_{i}\lambda_{v}(D_{i},P)<[(l-n+1)(n+1)+\epsilon]h(A,P).$
Proof.
This is a direct consequence of [7, Corollary of the Main Theorem(pg. 140)] applied with $Y_{1}=D_{1},...,Y_{q}=D_{q}$ and of the fact that if $D$ is an effective Cartier divisor, then the Seshadri constants therein simplify to
[…]
(see the discussion in [6, Page 126]) ∎
The next one is a new Schmidt subspace type of theorem for algebraic points.
Lemma 3.2.
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$ Let $S$ be a finite set of places of $k$ and $D_{1},...,D_{q}$ be effective ample Cartier divisors on $X$, defined over $k$, in $m$-subgeneral position with $m\geq n$. Let $A$ be an ample Cartier divisor on $X$ and let $d_{i}$ be positive integers such that $d_{i}A\equiv D_{i}$ for all $i$ and for all $v\in S$. Let $\epsilon>0$ and $\delta\in\mathbb{N}^{\geq 1}$. Then
$\sum_{v\in S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}\sum_{i=1}^{q}\dfrac{\lambda_{w}(D_{i},P)}{d_{i}}<[C(m,n,\delta)+\epsilon]h(A,P)+O(1)$
holds for all points $P\in X(\overline{k})\setminus\bigcup_{i}\text{Supp }D_{i}$ satisfying $[k(P):k]\leq\delta$, where $C(m,n,\delta)$ is a positive constant given in [11, Page 229].
Proof.
This is [11, Theorem 2], which improves the bound on [12, Theorem 5.2] together with [19, Lemma 3.4]. ∎
When $X$ is $\mathbb{P}^{2}$ and the $D_{j}$’s are lines in general position, $C(m,n,\delta)$ to 15/2 can be improved as follows
Lemma 3.3.
[12, Theorem 1.5] Let $S$ be a finite set of places of a number field $k$ let $L_{1},...,L_{q}\subset\mathbb{P}^{2}$ be lines over $k$, in general position. Let $\epsilon>0$. There there exists a finite union of lines $Z\subset X$ such that for all points $P\in\mathbb{P}^{2}(\overline{k})\setminus Z$ satisfying $[k(P):k]\leq 2$ we have
$\sum_{v\in S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}\sum_{i=1}^{q}{\lambda_{w}(L_{i},P)}<\left(\dfrac{15}{2}+\epsilon\right)h(P).$
4. Proofs and results
4.1. Sparsity of integral points in orbits
Here we present some new results in the context of the new diophantine approximation result of He and Ru.
Moreover, we prove results in semigroup dynamics. Namely, for $\phi_{1},...,\phi_{\ell}:X\rightarrow X$ not necessarily commuting, we consider $\mathcal{F}:=<\phi_{1},...,\phi_{\ell}>$ the semigroup generated by $\phi_{1},...,\phi_{\ell}$ with the composition of functions as its operation.
Theorem 4.1.
Let $X$ be a projective variety of dimension $n$, $S$ be a finite set of places, $\phi_{1},\ldots,\phi_{\ell}$ be endomorphisms of $X$, and $\mathcal{F}=\langle\phi_{1},\ldots,\phi_{\ell}\rangle$ all def. over a number field $k$. Suppose there exist $\psi_{1},\ldots,\psi_{m}\in\mathcal{F}$, a nontrivial effective divisor $D$ and $A$ an ample divisor of $X$ such that
  • (i)
    $\psi_{i}^{*}(D)=D_{i1}+D_{i2}+\cdots+D_{i,q_{i}}+D_{i}^{\prime}$ is a sum of effective divisors, where $D_{i1},\ldots,D_{i,q_{i}}$ are in $l_{i}$-subgeneral position for some $l_{i}\geq n$
  • (ii)
    There exist $c_{ij}\in\mathbb{Q}$ such that $\sum_{j=1}^{q_{i}}c_{ij}{D_{ij}}-(l_{i}-n+1)(n+1)A$ is big and $A-c_{ij}D_{ij}$ is $\mathbb{Q}$-nef for all $i,j$.
  • (iii)
    $\mathcal{F}\setminus\bigcup_{i=1}^{m}(\psi_{i}\circ\mathcal{F})$ is finite.
Then there exists $\epsilon>0$ such that for all $P\in X(k)$, the set
(4.1) $\big{\{}\phi(P):\phi\in\mathcal{F},\sum_{v\notin S}\lambda_{v}(D,\phi(P))\leq\epsilon h(A,\phi(P))\big{\}}$
is not Zariski dense.
Proof.
Suppose that $\phi(P)$ is in the set (4.1). By condition the (ii), we can suppose that $\phi=\psi_{i}\circ\eta$ for some $\eta\in\mathcal{F}$ and $1\leq i\leq m$. By functoriality of heights,
(4.2) $\begin{split}\sum_{v\notin S}&\lambda_{v}(D,\phi(P))\sim h(D,\psi_{i}(\eta(P)))-\sum_{v\in S}\lambda_{v}(D,\psi_{i}(\eta(P)))\\ &\sim h(\psi_{i}^{*}(D),\eta(P))-\sum_{v\in S}\lambda_{v}(\psi_{i}^{*}(D),\eta(P))\\ &\sim\sum_{v\notin S}\lambda_{v}(\psi_{i}^{*}(D),\eta(P))\\ &\geq\sum_{j=1}^{q_{i}}\sum_{v\notin S}\lambda_{v}(D_{ij},\eta(P))\end{split}$
where we use $\sim$ to indicate that the difference of both sides is a bounded function on Zariski-dense subset of $X(k)$. On the other hand, since $L_{i}:=\sum_{j=1}^{q_{i}}c_{ij}{D_{ij}}-(l_{i}-n+1)(n+1)A$ is big, there exists a proper Zariski-closed set $Z^{\prime}_{i}$ and a positive constant $c_{i}^{\prime}$ such that
(4.3) $h(L_{i},Q)\geq c_{i}^{\prime}h(A,Q)$
for all $Q\notin Z^{\prime}_{i}$. In addition, because of ampleness and
(4.4) $h(A,\psi_{i}(Q))\sim h(\psi_{i}^{*}(A),Q)$
for all $Q$ (functoriality), there exists $c_{i}^{\prime\prime}\geq 1$ not depending on $P$ such that
(4.5) $h(A,\psi_{i}(\eta(P)))\leq c_{i}^{\prime\prime}h(A,\eta(P)).$
Therefore, applying the above with $Q=\eta(P)$, combining with 4.3, 4.4, 4.5, and the condition in 4.1, we obtain
$h(-(l_{i}-n+1)(n+1)A,\eta(P))+\sum_{j=1}^{q_{i}}\sum_{v\in S}{\lambda_{v}(c_{ij}D_{ij},\eta(P))}$
$\sim h\left(-(l_{i}-n+1)(n+1)A+\sum_{j=1}^{q_{i}}c_{ij}{D_{ij}},\eta(P)\right)-\sum_{j=1}^{q_{i}}\sum_{v\notin S}{\lambda_{v}(c_{ij}D_{ij},\eta(P))}$
$\geq c_{i}^{\prime}h(A,\eta(P))-\left(\max_{j}c_{ij}\right)\sum_{v\notin S}\lambda_{v}(D,\phi(P))$
$\geq c_{i}^{\prime}h(A,\eta(P))-\epsilon\cdot\left(\max_{j}c_{ij}\right)h(A,\phi(P))$
$\geq c_{i}^{\prime}h(A,\eta(P))-\epsilon\max_{j}{c_{i}^{\prime\prime}}c_{ij}h(A,\eta(P)).$
Therefore, whenever
(4.6) $\epsilon<\displaystyle\min_{i}\left(\frac{c_{i}^{\prime}}{c_{i}^{\prime\prime}}\min_{j}1/c_{ij}\right),$
Lemma 3.1 applied $m$ times to each system of divisors $\{D_{ij}\}_{j=0,...,q_{i}}$ implies that there exists a union $Z_{1}$ of varieties such that $\eta(P)\in Z_{1}^{\prime}$, or it belongs to a set $Z_{2}$ of bounded height. Therefore, this argument shows that if $\phi(P)$ is in the set 4.1, then $\phi(P)=\psi_{i}(\eta(P))\in\psi_{i}(Z^{\prime}_{i})\cup\psi_{i}(Z_{1})\cup\psi_{i}(Z_{2})$, which concludes the proof. ∎
Corollary 4.2.
Let $X$ be a projective variety of dimension $n$, $S$ be a finite set of places, $\phi_{1},\ldots,\phi_{\ell}$ be endomorphisms of $X$, and $\mathcal{F}=\langle\phi_{1},\ldots,\phi_{\ell}\rangle$ all def. over a number field $k$. Suppose there exist $\psi_{1},\ldots,\psi_{m}\in\mathcal{F}$, a nontrivial effective divisor $D$ and $A$ an ample divisor of $X$ such that
  • (i)
    $\psi_{i}^{*}(D)=D_{i1}+D_{i2}+\cdots+D_{i,q_{i}}+D_{i}^{\prime}$ is a sum of effective divisors, $D_{i1},\ldots,D_{i,q_{i}}$ are in $l_{i}$-subgeneral position for some $l_{i}\geq n$, $\sum_{j=1}^{q_{i}}\frac{D_{ij}}{d_{ij}}-(l_{i}-n+1)(n+1)A$ is big for all $i$ and $D_{ij}\equiv d_{ij}A$ where $d_{ij}$ are positive integers for all $i,j$.
  • (ii)
    $\mathcal{F}\setminus\bigcup_{i=1}^{m}(\psi_{i}\circ\mathcal{F})$ is finite.
Then there exists $\epsilon>0$ such that for all $P\in X(k)$, the set:
$\big{\{}\phi(P):\phi\in\mathcal{F},\sum_{v\notin S}\lambda_{v}(D,\phi(P))\leq\epsilon h(A,\phi(P))\big{\}}$
is not Zariski dense
Proof.
This is a consequence of the previous theorem applied with $c_{ij}=1/d_{ij}.$
Corollary 4.3.
Let $X$ be a projective variety of dimension $n$, $S$ be a finite set of places and $\phi$ be an endomorphism of $X$ all def. over a number $k$. Suppose there exist $m\geq 0$, a nontrivial effective divisor $D$ and $A$ an ample divisor of $X$ such that
  • (i)
    $(\phi^{(m)})^{*}(D)=D_{1}+D_{2}+\cdots+D_{q}+D^{\prime}$ is a sum of effective divisors, $D_{1},\ldots,D_{q}$ are in $l$-subgeneral position for some $l\geq n$, $\sum_{j=1}^{q}\frac{D_{j}}{d_{j}}-(l-n+1)(n+1)A$ is big and $D_{j}\equiv d_{j}A$ $d_{j}$ are positive integers for all $j$.
Then there exists $\epsilon$ such that for all $P\in X(k)$, the set
$\big{\{}\phi^{(n)}(P):n\in\mathbb{N},\sum_{v\notin S}\lambda_{v}(D,\phi^{(n)}(P)\leq\epsilon h(A,\phi^{(n)}(P)\big{\}}$
is not Zariski dense.
4.2. Algebraic points of bounded degree
Here we prove some new statements on sparsity of integral points of bounded degree in orbits, that are also uniform in the sense that they don’t depend of the initial orbit points.
Theorem 4.4.
Let $X$ be a projective variety of dimension $n$, $S$ be a finite set of places, $\phi_{1},\ldots,\phi_{\ell}$ be endomorphisms of $X$, and $\mathcal{F}=\langle\phi_{1},\ldots,\phi_{\ell}\rangle$ all def. over a number $k$. Suppose there exist $\psi_{1},\ldots,\psi_{m}\in\mathcal{F}$, a nontrivial effective divisor $D$ and $A$ an ample divisor of $X$ such that
  • (i)
    $\psi_{i}^{*}(D)=D_{i1}+\cdots+D_{i,q_{i}}+D_{i}^{\prime}$ is a sum of effective ample divisors, where $D_{i1},\ldots,D_{i,q_{i}}$ are in $l_{i}$-subgeneral position with $l_{i}\geq n$
  • (ii)
    $\sum_{j=1}^{q_{i}}\frac{D_{ij}}{d_{ij}}-C(l_{i},n,\delta)A$ is big and $D_{ij}\equiv d_{ij}A$ $d_{ij}$ are positive integers for all $i,j$.
Then there exists $\epsilon>0$ such that the set
(4.7) $\left\{\phi(P):\phi\in\bigcup_{i=1}^{m}(\psi_{i}\circ\mathcal{F}),[k(P):k]\leq\delta,\sum_{v\notin S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}\lambda_{w}(D,\phi(P))\leq\epsilon h(A,\phi(P))\right\}$
is not Zariski dense.
Proof.
The proof follows the lines of the proof of Theorem 5.1, adapted to points in extensions of bounded degree. Suppose that $\phi(P)$ is in the set (4.7). Then $\phi=\psi_{i}\circ\eta$ for some $\eta\in\mathcal{F}$ and $1\leq i\leq m$. By functoriality of heights,
(4.8) $\begin{split}\sum_{v\notin S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}&\lambda_{v}(D,\phi(P))\sim h(D,\psi_{i}(\eta(P)))-\sum_{v\in S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}\lambda_{v}(D,\psi_{i}(\eta(P)))\\ &\sim h(\psi_{i}^{*}(D),\eta(P))-\sum_{v\in S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}\lambda_{v}(\psi_{i}^{*}(D),\eta(P))\\ &\sim\sum_{v\notin S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}\lambda_{v}(\psi_{i}^{*}(D),\eta(P))\\ &\geq\sum_{j=1}^{q_{i}}\sum_{v\notin S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}\lambda_{v}(D_{ij},\eta(P)).\end{split}$
On the other hand, since $L_{i}:=\sum_{j=1}^{q_{i}}\frac{D_{ij}}{d_{ij}}-C(l_{i},n,\delta)A$ is big, there exists a proper Zariski-closed set $Z^{\prime}_{i}$ and a positive constant $c_{i}^{\prime}$ such that
(4.9) $h(L_{i},Q)\geq c_{i}^{\prime}h(A,Q)$
for all $Q\notin Z_{i}$. In addition, because of ampleness and
(4.10) $h(A,\psi_{i}(Q))\sim h(\psi_{i}^{*}(A),Q)$
for all $Q$ (functoriality), there exists $c_{i}^{\prime\prime}\geq 1$ such that
(4.11) $h(A,\psi_{i}(\eta(P)))\leq c_{i}^{\prime\prime}h(A,\eta(P))$
for all $P$. Therefore, applying the above with $Q=\eta(P)$, combining with (4.9), (4.10), (4.11) and the condition in (4.7), we obtain
$h(-C(l_{i},n,\delta)A,\eta(P))+\sum_{j=0}^{q_{i}}\sum_{v\in S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}{\lambda_{w}\left(\frac{D_{ij}}{d_{ij}},\eta(P)\right)}$
$\sim h\left(-C(l_{i},n,\delta)A+\sum_{j=0}^{q_{i}}\frac{D_{ij}}{d_{ij}},\eta(P)\right)-\sum_{j=0}^{q_{i}}\sum_{v\notin S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}{\lambda_{w}\left(\frac{D_{ij}}{d_{ij}},\eta(P)\right)}$
$\geq c_{i}^{\prime}h(A,\eta(P))-\left(\max_{j}1/d_{ij}\right)\sum_{v\notin S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}\lambda_{w}(D,\phi(P))$
$\geq c_{i}^{\prime}h(A,\eta(P))-\epsilon\cdot\left(\max_{j}1/d_{ij}\right)h(A,\phi(P))$
$\geq c_{i}^{\prime}h(A,\eta(P))-\epsilon\max_{j}\frac{c_{i}^{\prime\prime}}{d_{ij}}h(A,\eta(P)).$
Therefore, whenever
(4.12) $\epsilon<\displaystyle\min_{i}\left(\frac{c_{i}^{\prime}}{c_{i}^{\prime\prime}}\min_{j}d_{ij}\right),$
Lemma 3.2 applied $m$ times to each system of divisors $\{D_{ij}\}_{j=1,...,q_{i}}$ implies that there exists a union $Z_{1}$ of varieties such that $\eta(P)\in Z_{1}$, or it belongs to a set $Z_{2}$ of bounded height. Therefore, this argument shows that if $\phi(P)$ is in the set (4.7), $\phi(P)=\psi_{i}(\eta(P))\in\psi_{i}(Z^{\prime}_{i})\cup\psi_{i}(Z_{1})\cup\psi_{i}(Z_{2})$, which concludes the proof. ∎
Theorem 4.5.
Let $\phi_{1},\ldots,\phi_{\ell}$ be endomorphisms of $\mathbb{P}^{2}$, and $\mathcal{F}=\langle\phi_{1},\ldots,\phi_{\ell}\rangle$ all def. over a number $k$. Suppose there exist $\psi_{1},\ldots,\psi_{m}\in\mathcal{F}$ and a nontrivial effective divisor $D$ such that
  • (i)
    $\psi_{i}^{*}(D)=L_{i1}+L_{i1}+\cdots+L_{i,q_{i}}+D_{i}^{\prime}$ a sum of effective divisors, $L_{i1},\ldots,L_{i,q_{i}}$ are lines in general position, and $q_{i}>15/2$ for all $i$.
Then there exists $\epsilon>0$ such that the set
$\left\{\phi(P):\phi\in\bigcup_{i=1}^{m}(\psi_{i}\circ\mathcal{F}),[k(P):k]\leq 2,\sum_{v\notin S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}\lambda_{w}(D,\phi(P))\leq\epsilon h(\phi(P))\right\}$
is not Zariski dense (here, h is the standard logarithmic height on $\mathbb{P}^{2}$).
Proof.
We can prove this similarly as in the proof of the previous theorem, adapting accordingly. In fact, we consider $X=\mathbb{P}^{2}$, $D_{ij}=L_{ij}$ to be lines in general position, and $A\sim O_{\mathbb{P}^{2}}(1)$ a hyperplane divisor representing a generator of the Picard group of $\mathbb{P}^{2}$, whose associated Weil height is the usual naive logarithmic height $h$ of $\mathbb{P}^{2}$ in the statement of the Theorem 5.4. In this case, the condition (ii) from Thm. 5.4 is satisfied with $d_{ij}=1$ and $q_{i}>15/2$, which is equivalent to having $\sum_{j=1}^{q_{i}}L_{ij}-(15/2)O_{\mathbb{P}^{2}}(1)$ big. With these data, the proof follows the same steps of the proof of Thm. 5.4 but replacing each $C(l_{i},n,\delta)$ with 15/2 and using Lemm 3.3 instead of Lemma 3.2. ∎
Corollary 4.6.
Let $\phi$ be an endomorphism of $\mathbb{P}^{2}$ both def. over a number $k$. Suppose there exist $m\geq 0$ and a nontrivial effective divisor $D$ such that
  • (i)
    $(\phi^{(m)})^{*}(D)=L_{1}+L_{2}+\cdots+L_{q}+D^{\prime}$ a sum of effective divisors, $L_{1},\ldots,L_{q}$ are lines in general position, and $q>15/2$
Then there exists $\epsilon$ such that the set
$\big{\{}\phi^{(n)}(P):n\in\mathbb{N}^{\geq m},[k(P):k]\leq 2,\sum_{v\notin S}\sum_{\begin{subarray}{c}w\mid v\\ w\in M_{k(P)}\end{subarray}}\lambda_{w}(D,\phi^{(n)}(P))\leq\epsilon h(\phi^{(n)}(P))\big{\}}$
is not Zariski dense.